a 3.00 kg object experiences a 16.0 N force pulling south, a 26.3 N force pulling east and a 21.0 N force pulling at 27 degrees S of E. The magnitude of the acceleration experienced by the object is?


1
Expert's answer
2021-11-18T15:09:38-0500

"F_x=26.3+21\cdot\cos27°=45.9\ (N)"


"F_y=16+21\cdot\sin27°=25.5\ (N)"


"F=\sqrt{45.9^2+25.5^2}=52.5\ (N)"


"a=F/m=52.5/3=17.5\ (m/s^2)" . Answer